# Random divergence-free drifts and the Onsager-Richardson threshold

### 12 Mar 2026 v1

@@ 2. General theorem and reduction to the Morse-Sard property
Both Theorem 1.2 and Theorem 1.3 are special cases of the following more general one. In order to state it we first recall the Hodge $\star$ operator on (the flat torus) $\mathbb{T}^{d}$, which is a canonical way of mapping $k$-forms into (dk)(d-k)-forms. For our purposes we are interested only in the case $k=1$. If $x_{1},\ldots,x_{d}$ are the canonical coordinates on $\mathbb{T}^{d}=\mathbb{R}^{d}/\mathbb{Z}^{d}$ the Hodge $\star$ maps the 1-form $\sum_{i=1}^{d}w_{i}dx_{i}$ to the d1d-1 form
@@
The next concept which we introduce is the “renormalization property” of DiPerna-Lions.
Definition 2.1.
Let $v$ be a bounded divergence-free vector field on $\mathbb{T}^{d}$ (not necessarily autonomous). A bounded weak solution $\theta$ of the transport equation
$\left\{\begin{array}[]{ll}\partial_{t}\theta+(v\cdot\nabla)\theta=0,\\ \\ \theta(\cdot,0)=\theta_{in},\end{array}\right$. (2.1)
is a bounded function $\theta$ which satisfies
$\int_{0}^{\infty}\int_{\mathbb{T}^{d}}\theta(\partial_{t}\varphi+v\cdot\nabla\varphi)\,dx\,dt=-\int_{\mathbb{T}^{d}}\theta_{in}(x)\varphi(x,0)\,dx$ (2.2)
The vector field $v$ is said to have the DiPerna-Lions renormalization property if, for every bounded initial data $\theta_{in}\in L^{\infty}(\mathbb{T}^{d})$, every bounded weak solution $\theta$ of (2.1), and every test function $\beta\in C^{1}(\mathbb{R})$, the scalar $\beta(\theta)$ is a weak solution of (2.1) in the sense of equation (2.2) with initial data $\beta(\theta_{in})$.
@@
Consider a probability measure $\mathbb{P}$ on the space of continuous, zero average, functions $\phi=(\phi_{1},\ldots,\phi_{d-1})\in C(\mathbb{T}^{d},\mathbb{R}^{d-1})$ satisfying the following two assumptions:
@@
    There is $\alpha>\frac{1}{3}$ such that $\phi\in C^{1,\alpha}(\mathbb{T}^{d};\mathbb{R}^{d-1})$ $\mathbb{P}$-almost surely;
@@
It is now easy to see that Theorem 1.2 and Theorem 1.3 are direct consequences of Theorem 2.2. In fact for $d=2$ the relation (2.4) is equivalent to $w=\nabla^{\perp}\phi_{1}$. Since $w$ is assumed, in Theorem 1.2, to be average-free, there is for each such $w$ a unique stream function $\phi_{1}:\mathbb{T}^{2}\to\mathbb{R}$ with zero average such that $w=\nabla^{\perp}\phi_{1}$: a probability measure as in Theorem 1.2 immediately translates into a probability measure as in Theorem 2.2. In the case of Theorem 1.3 a simple computation shows that (2.4) is in fact equivalent to $w=\nabla\phi_{1}\times\nabla\phi_{2}$. So far we have shown that the assumptions in Theorems 1.2 and 1.3 are in fact equivalent to those of Theorem 2.2 when $d=2$ and $3$, respectively.
The key to the proof of Theorem 2.2 is to show that $\phi$ satisfies the Morse-Sard property $\mathbb{P}$-almost surely and then use the deep result of Alberti, Bianchini, and Crippa in [2]. In particular, in Section 3 we will show the following.
Proposition 2.3.
Let $\mathbb{P}$ be a probability measure satisfying the same assumptions as in Theorem 2.2. Then $\mathbb{P}$-almost surely $\phi\in C^{1,\alpha}(\mathbb{T}^{d};\mathbb{R}^{d-1})$ has the Morse-Sard property, namely the set of its critical values $\{\phi(x):{\rm rank}\,(D\phi(x))\leq d-2\}$ is a (Lebesgue) null-set of $\mathbb{R}^{d-1}$.
With Proposition 2.3 at hand, we can use the results of [2] to conclude the renormalization property for any such $\phi$ which satisfies the Morse-Sard property. In the case of two-dimensional vector fields the statement follows immediately from Theorem 5.2 and the remark in Section 6.3 of [2]. In the case of higher-dimensional vector fields the statement follows from Section 6.2 and Section 6.3 in [2]. Note indeed that the components $\phi_{k}$ of the vector map $\phi$ satisfy $\nabla\phi_{k}\cdot w\equiv 0$, namely the condition required in [2, Section 6.2]. This orthogonality relation is easy to verify from the properties of the Hodge $\star$ operator. In fact for every vector field $z$ we have the relation
@@
3. Proof of the Morse-Sard property
@@
Consider a probability measure $\mathbb{P}$ as in Proposition 2.3, and a function $\phi$ satisfying conditions (a) and (b) from Theorem 2.2. Then the Hausdorff dimension of the critical set
$Z:=\{{\rm rank}\,(D\phi)\leq d-2\}$ (3.1)
is, $\mathbb{P}$-almost surely, at most $d-2\alpha$.
@@
Let $\phi\in C^{1,\alpha}(\mathbb{T}^{d},\mathbb{R}^{d-1})$ be a function such that the Hausdorff dimension of its critical set $Z$ is at most $d-2\alpha$. If $\alpha>\frac{1}{3}$, then the set $\phi(Z)$ is a (Lebesgue) null-set of $\mathbb{R}^{d-1}$.
Clearly Proposition 2.3 is a direct consequence of Lemmas 3.1 and 3.2.
Proof of Lemma 3.1.
@@
Step 1. We consider the Euclidean space $\mathbb{R}^{(d-1)\times d}$ of $(d-1)\times d$ real matrices and denote by $W$ the closed subset of matrices of rank at most d2d-2. If we denote by ${\rm dim}_{H}(E)$ the Hausdorff dimension of a set $E$ and introduce the set
$\mathcal{Y}:=\{\phi:{\rm dim}_{H}((D\phi)^{-1}(W))>d-2\alpha\}\,$,
then we can rephrase the claim of the theorem as $\mathbb{P}(\mathcal{Y})=0$. On the other hand $\mathcal{Y}$ is the countable union of
$\mathcal{Y}_{N}:=\{\phi:\|D\phi\|_{C^{1,\alpha}}\leq N\quad\mbox{and}\quad\mathcal{H}^{d-2\alpha+1/N}((D\phi)^{-1}(W\cap B_{N}))>0\}\,$, (3.2)
@@
Step 2. From now on we fix a positive natural number $N$ and aim at proving (3.3). First of all we recall that $W$ is a real-algebraic subset of $\mathbb{R}^{(d-1)\times d}$ of codimension $2$. Its intersection with the closed ball $B_{N}$ has, therefore, finite Minkowski (d(d1)2)(d(d-1)-2)-dimensional content. In other words there is a constant $C$ (depending on both $W$ and $N$ but not on $\delta$) such that, for every $\delta>0$,
$|\{A\in\mathbb{R}^{(d-1)\times d}:{\rm dist}\,(A,W\cap B_{N})\leq\delta\}|\leq C\delta^{2}\,.$
Consider now the cubical decomposition of $\mathbb{R}^{d(d-1)}$ in cubes $Q:=\delta M+[0,\delta)^{d(d-1)}$, for $M$ ranging in the lattice $\mathbb{Z}^{d(d-1)}$. The cubes are chosen so to be pairwise disjoint. We then let $\mathcal{Q}_{\delta}$ be the subcollection of such cubes which intersect the set $\{{\rm dist}\,(\cdot,W\cap B_{N})\leq\delta\}$. The cardinality of this collection is controlled by $C\delta^{-d(d-1)+2}$, where $C$ is another constant independent of $\delta$: this follows because the cubes are pairwise disjoint and their union is contained in $\{{\rm dist}\,(\cdot,W\cap B_{N})\leq\delta(1+\sqrt{d})\}$.
@@
$\mathbb{P}(\{\phi:{\rm dist}\,(D\phi(x),W\cap B_{N})\leq\delta\})\leq C\delta^{2}\,.$ (3.4)
Step 3. We next fix a natural number $k$, we subdivide $\mathbb{T}^{d}$ in $2^{-kd}$ closed nonoverlapping cubes $Q$ of sidelength $2^{-k}$, and for each $\phi\in C^{1,\alpha}$ we count the number $n(k,\phi)$ of cubes in the subdivision which intersect $(D\phi)^{-1}(W\cap B_{N})$. Let us enumerate the cubes of the subdivision as $Q_{i}$ and note that
$n(k,\phi)=\sum\mathbf{1}_{E_{i}}(\phi)$ (3.5)
@@
$n(k,\phi)\mathbf{1}_{\{\|\phi\|_{C^{1,\alpha}}\leq N\}}(\phi)\leq\sum_{i}\mathbf{1}_{F_{i}}(\phi)\,.$ (3.7)
By inequality (3.4), by choosing $\delta=N(\sqrt{d}\,2^{-k-1})^{\alpha}$ we have $\mathbb{P}(F_{i})\leq CN^{2}2^{-2k\alpha}$ and therefore, integrating (3.7) and by taking into account that there are $2^{kd}$ distinct cubes in the cubical decomposition of $\mathbb{T}^{d}$ considered above, we conclude immediately
$\int_{\|D\phi\|_{C^{1,\alpha}}\leq N}n(k,\phi)\,d\mathbb{P}(\phi)\leq\sum_{i}\mathbb{P}(F_{i})\leq CN^{2}2^{(d-2\alpha)k}\,.$ (3.8)
@@
$\mathcal{Y}_{N,k}:=\left\{\phi:\|D\phi\|_{C^{1,\alpha}}\leq N\quad\mbox{and}\quad n(k,\phi)\geq 2^{(d-2\alpha+1/(2N))k}\right\}\,$
@@
$\mathbb{P}(\mathcal{Y}_{N,k})\leq 2^{-(d-2\alpha+1/(2N))k}\mathbb{E}\big[n(k,\phi)\mathbf{1}_{\{\|\phi\|_{C^{1,\alpha}}\leq N\}(\phi)\big]}\leq CN^{2}2^{-k/(2N)}\,.$ (3.9)
@@
is a set of measure zero.
We now claim that $\mathcal{Y}_{N}\subset\mathcal{Y}^{\prime}_{N}$, which would conclude the proof. In fact, assume $\phi\in\mathcal{Y}_{N}$. If $\phi$ is in the complement of $\mathcal{Y}^{\prime}_{N}$ then necessarily there would be a sequence of natural numbers $k(i)\to\infty$ with the property that $\phi\not\in\mathcal{Y}_{N,k(i)}$ for every $i$. But, given the definition of $\mathcal{Y}$, $\|\phi\|_{C^{1,\alpha}}\leq N$. This means that, for every $i$, $n(k(i),\phi)<2^{(d-2\alpha+1/(2N))k(i)}$. Therefore we can cover $(D\phi)^{-1}(W\cap B_{N})$ with less then $2^{(d-2\alpha+1/(2N))k(i)}$ cubes of sidelength $2^{-k(i)}$. We then conclude, from the definition of the $\mathcal{H}^{\beta}_{\infty}$ premeasure, that
$\mathcal{H}^{d-2\alpha+N^{-1}}_{\infty}((D\phi)^{-1}(W\cap B_{N}))\leq\omega_{d-2\alpha+N^{-1}}2^{-k(i)/(2N)}\left(\frac{\sqrt{d}}{2}\right)^{d-2\alpha+N^{-1}}\,.$ (3.10)
But letting $i\uparrow\infty$ we would then conclude $\mathcal{H}^{d-2\alpha+N^{-1}}_{\infty}((D\phi)^{-1}(W\cap B_{N}))=0$. Since $\mathcal{H}^{\beta}_{\infty}$-null sets and $\mathcal{H}^{\beta}$-null sets coincide, we then would be in contradiction with the definition of $\mathcal{Y}_{N}$, which requires $\mathcal{H}^{d-2\alpha+N^{-1}}((D\phi)^{-1}(W\cap B_{N}))>0$. Therefore we have $\mathcal{Y}_{N}\subset\mathcal{Y}^{\prime}_{N}$, from which it follows that both $\mathbb{P}(\mathcal{Y}_{N})$ and $\mathbb{P}(\mathcal{Y})$ are equal to zero. ∎
Proof of Lemma 3.2.
Since $Z$ has Hausdorff dimension at most $d-2\alpha$ (as was proved in Lemma 3.1), for every positive $\delta$ we have $\mathcal{H}^{d-2\alpha+\delta}(Z)=0$. We fix here a $\delta$ such that
$3\alpha\geq 1+\delta\,.$ (3.11)
From the definition of Hausdorff measure it follows that for every positive $\varepsilon\leq 1$ there is a cover of $Z$ with sets $U_{k}$ such that
$\sum_{k}({\rm diam}\,(U_{k}))^{d-2\alpha+\delta}<\varepsilon\,.$ (3.12)
We then discard any $U_{k}$ which does not intersect $Z$ and for the remaining ones we pick a point $x_{i}\in U_{k}\cap Z$. Set $r_{i}:={\rm diam}\,(U_{i})$ and conclude that the closed balls $B_{r_{i}}(x_{i})$ cover $Z$ and satisfy (by rewriting equation (3.12))
$\sum_{k}r_{k}^{d-2\alpha+\delta}<\varepsilon\,.$ (3.13)
Next we recall that $\phi$ is Lipschitz and thus $\phi(U_{i})\subset B_{C_{0}r_{i}}(\phi(x_{i}))$, where $C_{0}={\rm Lip}\,(\phi)$. Moreover, the rank of the linear map $L_{i}:=D\phi(x_{i})$ is at most d2d-2, as $x_{i}\in Z$: this allows us to choose a (d2)(d-2)-dimensional vector space $V_{i}$ with the property that ${\rm Im}\,(L_{i})\subset V_{i}$.
Select then a unit normal vector $\nu_{i}$ which is orthogonal to $V_{i}$ and observe that, by Lagrange’s mean-value theorem, for every $y\in B_{r_{i}}(x_{i})$ there is a point $\xi$ on the segment $x_{i},y]$ such that
@@
In particular the orthogonal projection of $\phi(B_{r_{i}}(x_{i}))$ onto the orthogonal complement $V_{i}^{\perp}$ of $V_{i}$ is contained in a closed segment $\sigma_{i}$ of length $2[D\phi]_{0,\alpha}r_{i}^{1+\alpha}$ centered on the projection of $\phi(x_{i})$ onto $V_{i}^{\perp}$. If we denote by $z_{i}$ the projection of $\phi(x_{i})$ onto $V_{i}$, the arguments above prove the inclusion
$\phi(B_{r_{i}}(x_{i}))\subset(B_{C_{0}r_{i}}(z_{i})\cap V)+\sigma_{i}=:W_{i}\,.$

### 01 Apr 2026 v2

@@ 2.3. Higher dimensions
Theorem 1.4 is a special cases of the following more general one. In order to state it we first recall the Hodge $\star$ operator on (the flat torus) $\mathbb{T}^{d}$, which is a canonical way of mapping $k$-forms into (dk)(d-k)-forms. For our purposes we are interested only in the case $k=1$. If $x_{1},\ldots,x_{d}$ are the canonical coordinates on $\mathbb{T}^{d}=\mathbb{R}^{d}/\mathbb{Z}^{d}$ the Hodge $\star$ maps the 1-form $\sum_{i=1}^{d}w_{i}dx_{i}$ to the d1d-1 form
@@
Consider a probability measure $\mathbb{P}$ on the space of continuous functions $\phi=(\phi_{1},\ldots,\phi_{d-1})\in C(\mathbb{T}^{d},\mathbb{R}^{d-1})$ satisfying the following two assumptions:
@@
    There is $\alpha>\frac{1}{8}$ such that $\phi\in C^{1,\alpha}(\mathbb{T}^{d};\mathbb{R}^{d-1})$ $\mathbb{P}$-almost surely;
@@
The key to the proof of Theorem 2.4 is again to use the insight of Alberti, Bianchini, and Cripppa in [3]. First of all, note that the components $\phi_{k}$ of the vector map $\phi$ satisfy $\nabla\phi_{k}\cdot w\equiv 0$, namely one of the conditions required in [3, Section 6.2]. This orthogonality relation is easy to verify from the properties of the Hodge $\star$ operator. In fact for every vector field $z$ we have the relation
@@
Next, by Proposition 2.3 $\phi$ satisfies the weak Sard property almost surely. In this case, however, the weak Sard property alone is not enough to guarantee the renormalization property. In fact the authors in [3] need additionally the condition that “triods” are absent for a.e. level set of $\phi$ (for the precise definition of a triod we refer to [2, Section 2.3]) and in the subsequent paper [2] they show that this is not a consequence of the weak Sard property and might fail for $C^{1,\alpha}$ maps $\phi$. On the other hand, we can prove that this property does hold under the assumptions of Theorem 2.4.
Proposition 2.5.
Let $\mathbb{P}$ be a probability measure satisfying the same assumptions as in Theorem 2.4. Then $\mathbb{P}$-almost surely the map $\phi\in C^{1,\alpha}(\mathbb{T}^{d};\mathbb{R}^{d-1})$ satisfies the following property:
  • (NT)
    For a.e. $y\in\mathbb{R}^{d-1}$ the level set $\phi^{-1}(\{y\})$ contains no triods.
With Proposition 2.5 at hand, we can use the results of [3] to conclude the renormalization property for any such $\phi$ which satisfies the weak Morse-Sard property, cf. Section 6.2 and Section 6.3 in [3].
2.4. Absence of triods
@@
Let $d$ and $k\in\{0,\ldots,d-2\}$. Consider a probability measure $\mathbb{P}$ as in Proposition 2.5. Then the Hausdorff dimension of the set
$Y:=\{x\in\mathbb{T}^{d}:{\rm rank}\,(D\phi(x))\leq k\}$ (2.5)
is, $\mathbb{P}$-almost surely, at most $d-(d-k)(d-1-k)\alpha$.
Note that (dk)(d1k)(d-k)(d-1-k) is the codimension, in $\mathbb{R}^{d\times(d-1)}$ of the set of matrices with rank no larger than $k$.
@@
Let $d$ and $k$ be as in Lemma 2.6. Let $\phi\in C^{1,\alpha}(\mathbb{T}^{d},\mathbb{R}^{d-1})$ be a function such that the Hausdorff dimension of the set $Y$ in (2.5) is at most $\beta$. Then the set $\phi(Y)$ has Hausdorff dimension at most $\frac{\beta+\alpha k}{1+\alpha}$.
Before proving the lemma we show how Proposition 2.5 and Theorem 1.5 both follow from it.
Proof of Theorem 1.5.
Let $\mathbb{P}$ and $\alpha$ be as in the statement and choose k=d2k=d-2. In this case $Y$ in (2.5) is the set $Z$ of critical values of $\phi$, while $\phi(Z)$ is the set of critical points. But Lemma 2.6 implies that the dimension of $Z$ is, almost surely, no larger than $d-2\alpha$, while Lemma 2.7 implies then that $\phi(Z)$ has Hausdorff dimension no larger than $\frac{d-2\alpha+(d-2)\alpha}{1+\alpha}$. If the latter number is <d1<d-1, then $\phi(Z)$ is a (Lebesgue) null set of $\mathbb{R}^{d-1}$. On the other hand we have
$\frac{d-2\alpha+(d-2)\alpha}{1+\alpha}<d-1$
if and only if $\alpha>\frac{1}{3}$. ∎
Proof of Proposition 2.5.
Let $\mathbb{P}$ and $\alpha$ be as in Proposition 2.5. Consider k=d3k=d-3 and let $Y$ be as in (2.5). Note that $Y$ is closed and so its complement $U$ is open. According to [2, Proof of Lemma 2.16] the set
$\{y\in\mathbb{R}^{d-1}:\phi^{-1}(y)\cap U\mbox{ contains a triod}\}$
is a (Lebesgue) null set in $\mathbb{R}^{d-1}$. Therefore it suffices to show that
  • (A)
    $\mathbb{P}$-almost surely the set $\phi(Y)$ is a null set.
First observe that Lemma 2.6 gives that, $\mathbb{P}$-almost surely, the dimension of $Y$ is at most $d-6\alpha$. Hence Lemma 2.7 gives that, $\mathbb{P}$-almost surely, the dimension of $\phi(Y)$ is at most $\frac{d-6\alpha+(d-3)\alpha}{\alpha+1}$. As in the previous proof we just need the inequality
$\frac{d-6\alpha+(d-3)\alpha}{\alpha+1}<d-1$
to conclude (A). The above inequality is equivalent to $\alpha>\frac{1}{8}$, thus completing the proof. ∎
2.5. Proof of the probabilistic dimension bound (Lemma 2.6)
@@
Step 1. We consider the Euclidean space $\mathbb{R}^{(d-1)\times d}$ of $(d-1)\times d$ real matrices and denote by $W$ the closed subset of matrices of rank at most $k$. Denote by $c_{0}=(d-k)(d-1-k)$ its codimension. If we denote by ${\rm dim}_{H}(E)$ the Hausdorff dimension of a set $E$ and introduce the set
$\mathcal{Y}:=\{\phi:{\rm dim}_{H}((D\phi)^{-1}(W))>d-c_{0}\alpha\}\,$,
then we can rephrase the claim of the lemma as $\mathbb{P}(\mathcal{Y})=0$. On the other hand $\mathcal{Y}$ is the countable union of
$\mathcal{Y}_{N}:=\{\phi:\|D\phi\|_{C^{1,\alpha}}\leq N\quad\mbox{and}\quad\mathcal{H}^{d-c_{0}\alpha+1/N}((D\phi)^{-1}(W\cap B_{N}))>0\}\,$, (2.6)
@@
Step 2. From now on we fix a positive natural number $N$ and aim at proving (2.7). First of all we recall that $W$ is a real-algebraic subset of $\mathbb{R}^{(d-1)\times d}$ of codimension $c_{0}$. Its intersection with the closed ball $B_{N}$ has, therefore, finite Minkowski $(d(d-1)-c_{0})$-dimensional content. In other words there is a constant $C$ (depending on both $W$ and $N$ but not on $\delta$) such that, for every $\delta>0$,
$|\{A\in\mathbb{R}^{(d-1)\times d}:{\rm dist}\,(A,W\cap B_{N})\leq\delta\}|\leq C\delta^{c_{0}}\,.$
Consider now the cubical decomposition of $\mathbb{R}^{d(d-1)}$ in cubes $Q:=\delta M+[0,\delta)^{d(d-1)}$, for $M$ ranging in the lattice $\mathbb{Z}^{d(d-1)}$. The cubes are chosen so to be pairwise disjoint. We then let $\mathcal{Q}_{\delta}$ be the subcollection of such cubes which intersect the set $\{{\rm dist}\,(\cdot,W\cap B_{N})\leq\delta\}$. The cardinality of this collection is controlled by $C\delta^{-d(d-1)+c_{0}}$, where $C$ is another constant independent of $\delta$: this follows because the cubes are pairwise disjoint and their union is contained in $\{{\rm dist}\,(\cdot,W\cap B_{N})\leq\delta(1+\sqrt{d})\}$.
@@
$\mathbb{P}(\{\phi:{\rm dist}\,(D\phi(x),W\cap B_{N})\leq\delta\})\leq C\delta^{c_{0}}\,.$ (2.8)
Step 3. We next fix a natural number $k$, we subdivide $\mathbb{T}^{d}$ in $2^{kd}$ closed nonoverlapping cubes $Q$ of sidelength $2^{-k}$, and for each $\phi\in C^{1,\alpha}$ we count the number $n(k,\phi)$ of cubes in the subdivision which intersect $(D\phi)^{-1}(W\cap B_{N})$. Let us enumerate the cubes of the subdivision as $Q_{i}$ and note that
$n(k,\phi)=\sum_{i=1}^{2^{kd}}\mathbf{1}_{E_{i}}(\phi)$ (2.9)
@@
$n(k,\phi)\mathbf{1}_{\{\|\phi\|_{C^{1,\alpha}}\leq N\}}(\phi)\leq\sum_{i=1}^{2^{kd}}\mathbf{1}_{F_{i}}(\phi)\,.$ (2.11)
By inequality (2.8), by choosing $\delta=N(\sqrt{d}\,2^{-k-1})^{\alpha}$ we have $\mathbb{P}(F_{i})\leq CN^{c_{0}}2^{-c_{0}k\alpha}$ and therefore, integrating (2.11) and by taking into account that there are $2^{kd}$ distinct cubes in the cubical decomposition of $\mathbb{T}^{d}$ considered above, we conclude immediately
$\int_{\|D\phi\|_{C^{1,\alpha}}\leq N}n(k,\phi)\,d\mathbb{P}(\phi)\leq\sum_{i=1}^{2^{kd}}\mathbb{P}(F_{i})\leq CN^{c_{0}}2^{(d-c_{0}\alpha)k}\,.$ (2.12)
@@
$\mathcal{Y}_{N,k}:=\left\{\phi:\|D\phi\|_{C^{1,\alpha}}\leq N\quad\mbox{and}\quad n(k,\phi)\geq 2^{(d-c_{0}\alpha+1/(2N))k}\right\}\,$
@@
$\mathbb{P}(\mathcal{Y}_{N,k})\leq 2^{-(d-c_{0}\alpha+1/(2N))k}\,\mathbb{E}\big[n(k,\phi)\mathbf{1}_{\{\|\phi\|_{C^{1,\alpha}}\leq N\}(\phi)\big]}\leq CN^{c_{0}}2^{-k/(2N)}\,.$ (2.13)
@@
is a set of probability zero.
We now claim that $\mathcal{Y}_{N}\subset\mathcal{Y}^{\prime}_{N}$, which would conclude the proof. In fact, assume $\phi\in\mathcal{Y}_{N}$. If $\phi$ is in the complement of $\mathcal{Y}^{\prime}_{N}$ then necessarily there would be a sequence of natural numbers $k(i)\to\infty$ with the property that $\phi\not\in\mathcal{Y}_{N,k(i)}$ for every $i$. But, given the definition of $\mathcal{Y}_{N}$, $\|\phi\|_{C^{1,\alpha}}\leq N$. This means that, for every $i$, $n(k(i),\phi)<2^{(d-c_{0}\alpha+1/(2N))k(i)}$. Therefore we can cover $(D\phi)^{-1}(W\cap B_{N})$ with less then $2^{(d-c_{0}\alpha+1/(2N))k(i)}$ cubes of sidelength $2^{-k(i)}$. We then conclude, from the definition of the $\mathcal{H}^{\beta}_{\infty}$ premeasure, that
$\mathcal{H}^{d-c_{0}\alpha+N^{-1}}_{\infty}((D\phi)^{-1}(W\cap B_{N}))\leq\omega_{d-c_{0}\alpha+N^{-1}}2^{-k(i)/(2N)}\left(\frac{\sqrt{d}}{2}\right)^{d-c_{0}\alpha+N^{-1}}\,.$ (2.14)
But letting $i\uparrow\infty$ we would then conclude $\mathcal{H}^{d-c_{0}\alpha+N^{-1}}_{\infty}((D\phi)^{-1}(W\cap B_{N}))=0$. Since $\mathcal{H}^{\beta}_{\infty}$-null sets and $\mathcal{H}^{\beta}$-null sets coincide, we then would be in contradiction with the definition of $\mathcal{Y}_{N}$, which requires $\mathcal{H}^{d-c_{0}\alpha+N^{-1}}((D\phi)^{-1}(W\cap B_{N}))>0$. Therefore we have $\mathcal{Y}_{N}\subset\mathcal{Y}^{\prime}_{N}$, from which it follows that both $\mathbb{P}(\mathcal{Y}_{N})$ and $\mathbb{P}(\mathcal{Y})$ are equal to zero.
2.6. Proof of Lemma 2.7
From the definition of Hausdorff measure it follows that for every positive $\varepsilon\leq 1$ and $\delta>0$ there is a cover of $Y$ with sets $U_{k}$ such that
$\sum_{k}({\rm diam}\,(U_{k}))^{\beta+\delta}<\varepsilon\,.$ (2.15)
We then discard any $U_{k}$ which does not intersect $Y$ and for the remaining ones we pick a point $x_{i}\in U_{k}\cap Y$. Set $r_{i}:={\rm diam}\,(U_{i})$ and conclude that the closed balls $B_{r_{i}}(x_{i})$ cover $Y$ and satisfy (by rewriting equation (2.15))
$\sum_{k}r_{k}^{\beta+\delta}<\varepsilon\,.$ (2.16)
Next we recall that $\phi$ is Lipschitz and thus $\phi(U_{i})\subset B_{C_{0}r_{i}}(\phi(x_{i}))$, where $C_{0}={\rm Lip}\,(\phi)$. Moreover, the rank of the linear map $L_{i}:=D\phi(x_{i})$ is at most $k$, as $x_{i}\in Y$: this allows us to choose a $k$-dimensional vector space $V_{i}$ with the property that ${\rm Im}\,(L_{i})\subset V_{i}$.
Select then any unit normal vector $\nu_{i}$ which is orthogonal to $V_{i}$ and observe that, by Lagrange’s mean-value theorem, for every $y\in B_{r_{i}}(x_{i})$ there is a point $\xi$ on the segment $x_{i},y]$ such that
@@
In particular the orthogonal projection of $\phi(B_{r_{i}}(x_{i}))$ onto the orthogonal complement $V_{i}^{\perp}$ of $V_{i}$ is contained in a closed (dk1)(d-k-1)-dimensional ball $\sigma_{i}$ of diameter bounded $2[D\phi]_{0,\alpha}r_{i}^{1+\alpha}$ centered on the projection of $\phi(x_{i})$ onto $V_{i}^{\perp}$. If we denote by $z_{i}$ the projection of $\phi(x_{i})$ onto $V_{i}$, the arguments above prove the inclusion
$\phi(B_{r_{i}}(x_{i}))\subset(B_{C_{0}r_{i}}(z_{i})\cap V_{i})+\sigma_{i}=:W_{i}\,.$